# A text-to-tabular approach to generate synthetic patient data using LLMs

### v1

@@ Abstract
Access to large-scale high-quality healthcare databases is key to accelerate medical research and make insightful discoveries about diseases. However, access to such data is often limited by patient privacy concerns, data sharing restrictions and high costs. To overcome these limitations, synthetic patient data has emerged as an alternative. However, synthetic data generation (SDG) methods typically rely on machine learning (ML) models trained on original data, leading back to the data scarcity problem. We propose an approach to generate synthetic tabular patient data that does not require access to the original data, but only a description of the desired database. We leverage prior medical knowledge and in-context learning capabilities of large language models (LLMs) to generate realistic patient data, even in a low-resource setting. We quantitatively evaluate our approach against state-of-the-art SDG models, using fidelity, privacy, and utility metrics. Our results show that while LLMs may not match the performance of state-of-the-art models trained on the original data, they effectively generate realistic patient data with well-preserved clinical correlations. An ablation study highlights key elements of our prompt contributing to high-quality synthetic patient data generation. This approach, which is easy to use and does not require original data or advanced ML skills, is particularly valuable for quickly generating custom-designed patient data, supporting project implementation and providing educational resources.
Keywords Synthetic data generation  $\cdot$ LLMs
Data and Code Availability
@@
This study uses data from the Parkinson’s Progression Markers Initiative (PPMI) database (RRID:SCR 006431) and from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (RRID:SCR 003007). For up-to-date information on these studies, visit www.ppmi-info.org and www.adni-info.org. The code used in this study is available at https://github.com/quinten-health-os/synth-data-gen-from-text.
1 Introduction
@@
Accelerating medical research using machine learning (ML) requires the availability of large, comprehensive, and complete patient data. In fact, to extract clinically relevant insights, ML models typically require access to extensive datasets that accurately represent the target population, maintain balanced class distributions, and without too many missing values. However, concerns regarding patient privacy, cost, and availability often limit access to such valuable data. Synthetic data generation (SDG) methods that aim at generating data that mimic real world data (RWD) have emerged as a promising solution to bypass these limits [Gonzales et al., 2023]. Indeed, synthetic data which maintain statistical properties of the original data while preserving patient privacy, can be used for a variety of analyzes to accelerate medical research. These include privacy-preserved data sharing [Yale et al., 2020, Hernandez et al., 2022, Eckardt et al., 2023], data augmentation [Perez and Wang, 2017, Seedat et al., 2024, Nakada et al., 2024] and cohort enrichment [Pauley et al., 2023]. In addition, SDG could accelerate and reduce the costs of randomized clinical trials (RCTs) by enhancing or replacing control arms with digital twins [L’Hostis et al., 2023, Alam et al., 2024].
@@
In this study, we present an innovative method for generating synthetic patient data using large language models (LLMs). Since LLMs are trained on biomedical literature databases such as PubMed, we hypothesize that these models could extract meaningful relationships between diseases, clinical symptoms, and patient characteristics, which is certainly key for generating high-quality synthetic patient data. As originally proposed by Borisov et al. [2023], we use an LLM directly as the SDG model. However, unlike previous tabular-to-tabular approaches relying on a learning phase on the original data and a generation phase of synthetic tabular data, we introduce a novel text-to-tabular approach requiring no original data and no pre-training or fine-tuning of the model. This method involves generating synthetic tabular data directly from a textual description of the desired database. The main motivation is to circumvent the need for inaccessible original patient-level data, and the high computational cost of training or fine-tuning SDG models.
@@
In this work, we present a framework for generating synthetic tabular patient data. Our contributions are the following:
  • We propose a text-to-tabular SDG approach using a frozen LLM (neither fine-tuned nor pre-trained) as the SDG model and only a database description as input. Unlike most generative models for tabular data, our approach does not require access to the original data. This makes it particularly useful in situations where the original data is difficult or timely to obtain or subject to privacy concerns.
@@
2 Related Works
Following advances in generative AI for images and text, attention has recently shifted towards tabular data. Various tabular-to-tabular SDG methods, involving a training phase of a model on the original data, have shown promising results in generating realistic RWD [Borisov et al., 2024].
@@ 2 Related Works
SDG can be effectively achieved using probabilistic graphical models such as Bayesian networks [Young et al., 2009, Zhang et al., ] and Gaussian Copulas (GCs), which model the joint distribution of variables from their marginal distributions [Patki et al., 2016, Kamthe et al., 2021].
@@
have demonstrated remarkable capabilities in image generation [Goodfellow et al., 2014], and have been successfully extended to tabular data [Borisov et al., 2024]. Among the recent advances, CTGAN stands out for its effectiveness in synthesizing mixed-type data i.e. data with a mix of both continuous and discrete variables, a long-standing challenge in tabular data generation [Xu et al., 2019]. CTGAN introduced a conditional vector in the generation process to handle unbalanced datasets and mode-specific normalization to effectively represent distributions with multiple modes. Building on this, [Zhao et al., 2021] introduced CTAB-GAN, which incorporates additional downstream losses and a novel encoding technique for the conditional vector, allowing efficient modeling of unbalanced or skewed data and mixed-type variables. Other tabular GANs, such as DP-GAN [Xie et al., 2018], Pate-GAN [Jordon et al., 2018] and CTAB-GAN+ [Zhao et al., 2022], also introduced specific losses and training techniques to ensure the privacy of the original data.
@@
Alongside CTGAN, Xu et al. [2019] proposed TVAE, which outperformed their suggested GAN approach. Diffusion models, which have gained significant attention in the computer vision domain [Austin et al., 2021], have also inspired approaches to generate tabular data. Notable examples include TABDDPM, which has shown remarkable performance in modeling heterogeneous features [Kotelnikov et al., 2022], and diffusion flow-based gradient boosted trees [Jolicoeur-Martineau et al., 2024].
@@
The application of LLMs, which have demonstrated significant proficiency in the textual domain, has inspired many to explore their use in the generation of realistic synthetic tabular data. The results obtained have been unexpectedly promising, outperforming state-of-the-art GANs in benchmark tabular datasets [Borisov et al., 2023], indicating the potential of LLMs in this area. Currently, most SDG approaches using LLMs involve fine-tuning the models on the original data or custom prompts including the original data. Borisov et al. [2023] pioneered the use of LLMs for SDG of tabular data, introducing GReaT, a method that uses pre-trained LLMs and that significantly outperforms GANs in synthesizing high-quality tabular data. Building on this foundation, several LLM-based SDG models with new features were developed: ReaLTAbFormer [Solatorio and Dupriez, 2023], TabuLa [Zhao et al., 2023] and DP-LLMTGen [Tran and Xiong, 2024]. Additionally, Seedat et al. [2024] proposed CLLM, a curated LLM for data augmentation in small datasets, utilizing a frozen LLM and an add-on curation filter to generate high-quality and diverse new samples using the original data. In the latest research has been proposed MALLM-GAN, a framework inspired by the architecture of GANs using a frozen LLM within both the generator and optimizer to produce synthetic data [Ling et al., 2024]. Moreover, Xu et al. [2024] showed that random order permutation of variables during fine-tuning, as proposed by Borisov et al. [2023], does not allow an accurate estimation of the conditional mixture of distributions and that permutations corresponding to learned functional dependencies between variables are more appropriate.
@@
All the works cited above require access to original data. A notable work introducing a LLM not requiring access to some is LLM-VP, a virtual patient generator that autorefines prompts through iterative interaction with an LLM and clinicians [Cook, 2024]. However, although relevant to SDG, it is primarily designed to simulate real-life clinical scenarios for educational purposes, rather than for SDG in scenarios where patient data is explicitly unavailable.
3 Materials and Methods
3.1 Data
3.1.1 PPMI
@@
PD is a complex, heterogeneous, age-related neurodegenerative disease that is often described by a combination of core diagnostic symptoms, including bradykinesia, rigidity, tremor and postural instability [Politis et al., 2010]. PPMI is a global observational study to establish biomarker-defined cohorts and identify clinical, imaging, genetic, and biospecimen markers of PD progression. The PPMI Clinical database includes data from PD, prodromal (at risk of PD or with another form of parkinsonism), and healthy participants, who were recruited at nearly 50 international sites since 2010. The database contains a wide range of data, including demographics, medical history, clinical assessments (e.g., the Movement Disorder Society - Unified Parkinson Disease Rating Scale (MDS-UPDRS) questionnaire, the Montreal Cognitive Assessment (MoCA) score), medication, biospecimens and proteomics, genetics, and imaging results. We used two curated datacuts obtained from PPMI upon request after approval from the PPMI Data Access Committee: #PPMI2024 (v.20240129) comprising 3,096 patients, which was used to evaluate our approach, datacut #PPMI2020 (v.2020420) comprising 683 patients, which was used to conduct our ablation study.
3.1.2 ADNI
@@
AD is a neurodegenerative disease that leads to noticeable memory loss, cognitive decline in thinking and reasoning abilities, and significant disruptions in social or occupational functioning [Castellani et al., 2010]. The ADNI was launched in 2003 as a public-private partnership, led by Principal Investigator Michael W. Weiner, MD. The primary goal of ADNI has been to test whether serial magnetic resonance imaging, positron emission tomography, other biological markers, and clinical and neuropsychological assessment can be combined to measure the progression of mild cognitive impairment and early AD. More generally, ADNI data is a comprehensive and widely used collection of longitudinal clinical, imaging, genetic, and other biomarker data. It encompasses various data types, including structural, functional, and molecular brain imaging, biofluid biomarkers, cognitive assessments, genetic data, and demographic information. We used a datacut #ADNI (ADNIMERGED_15Oct2024) obtained from ADNI upon request after approval from the ADNI Data Access Committee.
@@
We selected PPMI and ADNI to evaluate our approach due to their extensive use in research, with over 400 scientific publications based on its data for PPMI and 5,500 for ADNI. Of note, we intentionally selected medical datasets with a restricted access to address concerns regarding the use of LLMs trained on open-source data, not always known, and thus at risk of memorization by the LLM [Bordt et al., 2024].
3.2 Data preparation
@@
To ensure the selection of a representative population of PD and AD, we included patients diagnosed with confirmed PD and AD and captured their clinical characteristics at the time of diagnosis. We then selected a subset of variables characteristic of the clinical picture of PD (and respectively, AD) and well-represented in PPMI (for PD) and ADNI (for AD) datasets, as indicated by a low missingness rate (Table 6 and Table 7 in Appendix B). For comparison of our approach with tabular-to-tabular models that do not handle missing values, we excluded patients with missing values. This affected the final cohort size with a loss of 19% of patients in #PPMI2024, 3% patients in #PPMI2020, and 12% patients in #ADNI (Table 4 and Table 5 in Appendix A). To demonstrate the high practicality of LLMs, no preprocessing was performed on column names or values. Notably, we did not perform categorical value encoding, normalization, or outlier removal. Encoding of categorical values into their definition (e.g., "Male" or "Female" instead of 1 or 0) would have impacted our results, as LLMs are designed to handle text and not numerical values. However, for simplicity and to ease the replication or pursue of our work, we pautoreferred to add the categorical values and the associated mapping (e.g., 1 $\rightarrow$ "Male", 0 $\rightarrow$ "Female") in the data specifications of the prompt rather than in the preprocessing step.
3.3 Text-to-tabular synthetic data generation
Approach
@@ 3.3 Text-to-tabular synthetic data generation
The text - Prompt design
@@
Prior knowledge:
The prompt included prior knowledge of a medical database and a disease, that is, a comprehensive description of the medical database, detailing its name and nature, as well as the inclusion/exclusion criteria relevant to the population (e.g., disease, demographic criteria).
Instructions:
We configured the LLM to minimize verbosity by instructing it not to comment on, repeat, or answer the question. Given data specifications included detailed column names, medical definitions for each variable, the type of each variable (e.g., float, integer), and the encoding of categorical variables. The output format was specified to be in JSON, with a predetermined number of rows $n$, and no missing values.
Context:
To improve the realism of the synthetic data generated by the LLM, we leveraged its few-shot learning capabilities by including a limited number of fictitious patient examples. These examples can be derived from aggregated data found in the literature or by consulting medical experts. When using real patient records for generating examples, patient privacy must be considered. In our case, we added a single example created using the average values of the selected variables extracted from publications on #PPMI2020 [Marek et al., 2018] and on the ADNI dataset [Manning et al., 2017, Kikuchi et al., 2022].
The prompts designed to generate our synthetic PPMI and ADNI datasets are available upon request.
The tabular SDG process
@@
Ordering of columns being irrelevant in tabular datasets, we randomly permuted the columns at each generation, as proposed by Borisov et al. [2023].
3.4 Evaluation framework of synthetic patient data generation
To obtain a comprehensive assessment, we evaluated both fidelity, i.e. statistical similarity, privacy preservation and utility, i.e. added value of the synthetic data to a specific ML task or healthcare use case, using several widely used metrics in the SDG field or proposed by the Synthetic Data Vault in the Python library SDMetrics [Dat, 2024].
Evaluation framework
We compared our approach with several baseline tabular-to-tabular SDG models that require access to the original data during training: CTGAN [Xu et al., 2019], TVAE [Xu et al., 2019], and Gaussian copula (GC)[Kamthe et al., 2021]. CTGAN and TVAE were both trained for 300 epochs and using a batch size of 50. We chose a smaller batch size than the one proposed by Xu et al. [2019] because of the smaller training database.

### v2

# Margaux Törnqvist Quinten Health Paris, France

@@ Abstract
Access to large-scale high-quality healthcare databases is key to accelerate medical research and make insightful discoveries about diseases. However, access to such data is often limited by patient privacy concerns, data sharing restrictions and high costs. To overcome these limitations, synthetic patient data has emerged as an alternative. However, synthetic data generation (SDG) methods typically rely on machine learning (ML) models trained on original data, leading to the data scarcity problem. We propose an approach to generate synthetic tabular patient data that does not require access to the original data, but only a description of the desired database. We harness prior medical knowledge and the in-context learning capabilities of large language models (LLMs) to perform zero-shot generation of realistic patient data, even in low-resource settings. We quantitatively evaluate our approach against state-of-the-art SDG Models, using fidelity, privacy, and utility metrics. Our results show that while LLMs may not match the performance of state-of-the-art models trained on the original data, they effectively generate realistic patient data with well-preserved clinical correlations. An ablation study highlights key elements of our prompt that contribute to the generation of high-quality synthetic patient data. This approach, which is easy to use and does not require original data or advanced ML skills, is particularly valuable for quickly generating custom-designed patient data, supporting project implementation, and providing educational resources.
Index Terms:
GenAI, Synthetic data generation, LLM
@@
This study uses data from the Parkinson’s Progression Markers Initiative (PPMI) database (RRID:SCR 006431) and from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (RRID:SCR 003007). For up-to-date information on these studies, visit www.ppmi-info.org and www.adni-info.org. The code used in this study and Supplementary Materials are available at https://github.com/quinten-health-os/synth-data-gen-from-text.
I Introduction
@@
Accelerating medical research using machine learning (ML) requires the availability of large, comprehensive, and complete patient data. In fact, to extract clinically relevant insights, ML models typically require access to extensive datasets that accurately represent the target population, maintain balanced class distributions, and without too many missing values. However, concerns regarding patient privacy, cost, and availability often limit access to such valuable data. Synthetic data generation (SDG) methods that aim at generating data that mimic real world data (RWD) have emerged as a promising solution to bypass these limits [1]. Indeed, synthetic data which maintain statistical properties of the original data while preserving patient privacy, can be used for a variety of analyzes to accelerate medical research. These include privacy-preserved data sharing [2, 3, 4], data augmentation [5, 6, 7] and cohort enrichment [8]. In addition, SDG could accelerate and reduce the costs of randomized clinical trials (RCTs) by enhancing or replacing control arms with digital twins [9, 10].
@@
In this study, we present an innovative method for generating synthetic patient data using large language models (LLMs). Since LLMs are partially trained on biomedical corpora—including publicly available abstracts from databases like PubMed, we hypothesize that these models could extract meaningful relationships between diseases, clinical symptoms, and patient characteristics, which is certainly key for generating high-quality synthetic patient data. As originally proposed by [11], we use a pre-trained LLM directly as the SDG Model. However, unlike previous tabular-to-tabular approaches relying on a learning phase on the original data and a generation phase of synthetic tabular data, we introduce a novel text-to-tabular approach requiring no original data and no fine-tuning of the model. This method involves generating synthetic tabular data directly from a textual description of the desired database. The main motivation is to circumvent the need for inaccessible original patient-level data, and the high computational cost of training or fine-tuning SDG Models.
@@
In this work, we present a framework for generating synthetic tabular patient data. Our contributions are as follows.
  • We propose a text-to-tabular SDG approach using a frozen LLM (pre-trained but not fine-tuned) as the SDG Model conditioned solely on a database description. Unlike most generative models for tabular data, our approach does not require access to the original data. This makes it particularly useful in situations where the original data is difficult or timely to obtain or subject to privacy concerns.
@@
II Related Works
Following advances in generative AI for images and text, attention has recently shifted towards tabular data. Various tabular-to-tabular SDG methods, involving a training phase of a model on the original data, have shown promising results in generating realistic RWD [12].
@@ II Related Works
SDG can be effectively achieved using probabilistic graphical models such as Bayesian networks [13, 14] and Gaussian Copulas (GCs), which model the joint distribution of variables from their marginal distributions [15, 16].
@@
have demonstrated remarkable capabilities in image generation [17], and have been successfully extended to tabular data [12]. Among recent advances, CTGAN stands out for its effectiveness in synthesizing mixed-type data, that is, data with a mix of continuous and discrete variables, a long-standing challenge in tabular data generation [18]. CTGAN introduced a conditional vector in the generation process to handle unbalanced datasets and mode-specific normalization to effectively represent distributions with multiple modes. Building on this, [19] introduced CTAB-GAN, which incorporates additional downstream losses and a novel encoding technique for the conditional vector, allowing efficient modeling of unbalanced or skewed data and mixed-type variables. Other tabular GANs, such as DP-GAN [20], Pate-GAN [21] and CTAB-GAN+ [22], also introduced specific losses and training techniques to ensure the privacy of the original data.
@@
Alongside CTGAN, [18] proposed TVAE, which outperformed their suggested GAN approach. Diffusion models, which have gained significant attention in the computer vision domain [23], have also inspired approaches to generate tabular data. Notable examples include TABDDPM, which has shown remarkable performance in modeling heterogeneous features [24], and diffusion flow-based gradient boosted trees [25].
@@
The application of LLMs, which have demonstrated significant proficiency in the textual domain, has inspired many to explore their use in the generation of realistic synthetic tabular data. The results obtained have been unexpectedly promising, outperforming state-of-the-art GANs in benchmark tabular datasets [11], indicating the potential of LLMs in this area. Currently, most SDG approaches that use LLMs involve fine-tuning the models on the original data or custom prompts that include the original data. [11] pioneered the use of LLMs for the SDG of tabular data, introducing GReaT, a method that uses pre-trained LLMs and that significantly outperforms GANs in synthesizing high-quality tabular data. Building on this foundation, several LLM-based SDG Models with new features were developed: ReaLTAbFormer [26], TabuLa [27] and DP-LLMTGen [28]. Additionally, [6] proposed CLLM, a curated LLM for data augmentation in small datasets, utilizing a frozen LLM and an add-on curation filter to generate high-quality and diverse new samples using the original data. In the latest research has been proposed MALLM-GAN, a framework inspired by the architecture of GANs using a frozen LLM within both the generator and optimizer to produce synthetic data [29]. Moreover, [30] showed that the random order permutation of variables during fine-tuning, as proposed by [11], does not allow an accurate estimation of the conditional mixture of distributions and that permutations corresponding to learned functional dependencies between variables are more appropriate.
@@
All the works cited above require access to the original data. A notable work introducing a LLM that does not require access to some is LLM-VP, a virtual patient generator that refines prompts through iterative interaction with an LLM and clinicians [31]. However, although relevant to SDG, it is primarily designed to simulate real-life clinical scenarios for educational purposes, rather than for SDG in scenarios where patient data is explicitly unavailable.
III Materials and Methods
III-A Data
III-A1 PPMI
@@
Parkinson’s disease (PD) is a complex, heterogeneous, age-related neurodegenerative disease that is often described by a combination of core diagnostic symptoms, including bradykinesia, rigidity, tremor and postural instability [32]. PPMI is a global observational study to establish biomarker-defined cohorts and identify clinical, imaging, genetic, and biospecimen markers of PD progression. The PPMI Clinical database includes data from PD, prodromal (at risk of PD or with another form of parkinsonism), and healthy participants, who were recruited at nearly 50 international sites since 2010. The database contains a wide range of data, including demographics, medical history, clinical assessments (e.g., the Movement Disorder Society - Unified Parkinson Disease Rating Scale (MDS-UPDRS) questionnaire, the Montreal Cognitive Assessment (MoCA) score), medication, biospecimens and proteomics, genetics, and imaging results. We used two curated datacuts obtained from PPMI upon request after approval from the PPMI Data Access Committee: #PPMI2024 (v.20240129) comprising 3,096 patients, which was used to evaluate our approach, datacut #PPMI2020 (v.2020420) comprising 683 patients, which was used to conduct our ablation study.
III-A2 ADNI
@@
AD is a neurodegenerative disease that leads to noticeable memory loss, cognitive decline in thinking and reasoning abilities, and significant disruptions in social or occupational functioning [33]. The ADNI was launched in 2003 as a public-private partnership, led by Principal Investigator Michael W. Weiner, MD. The primary goal of ADNI has been to test whether serial magnetic resonance imaging, positron emission tomography, other biological markers, and clinical and neuropsychological assessment can be combined to measure the progression of mild cognitive impairment and early AD. More generally, ADNI data is a comprehensive and widely used collection of longitudinal clinical, imaging, genetic, and other biomarker data. It encompasses various data types, including structural, functional, and molecular brain imaging, biofluid biomarkers, cognitive assessments, genetic data, and demographic information. We used a datacut #ADNI (ADNIMERGED_15Oct2024) obtained from ADNI upon request after approval from the ADNI Data Access Committee.
@@
We selected PPMI and ADNI to evaluate our approach due to their extensive use in research, with over 400 scientific publications based on its data for PPMI and 5,500 for ADNI. Of note, we intentionally selected medical datasets with a restricted access to address concerns regarding the use of LLMs trained on open-source data, not always known, and thus at risk of memorization by the LLM [34].
III-B Data preparation
@@
To ensure the selection of a representative population of PD and AD, we included patients diagnosed with confirmed PD and AD and captured their clinical characteristics at the time of diagnosis. We then selected a subset of variables characteristic of the clinical picture of PD (and respectively, AD) and well-represented in PPMI (for PD) and ADNI (for AD) datasets, as indicated by a low missingness rate (Table VII and Table VIII in Supplementary Materials). For comparison of our approach with tabular-to-tabular models that do not handle missing values, we excluded patients with missing values. This affected the final cohort size with a loss of 19% of patients in #PPMI2024, 3% patients in #PPMI2020, and 12% patients in #ADNI (Table III and Table IV in Appendix A). To demonstrate the high practicality of LLMs, no preprocessing was performed on column names or values. Notably, we did not perform categorical value encoding, normalization, or outlier removal. Encoding of categorical values into their definition (e.g., ”Male” or ”Female” instead of 1 or 0) would have impacted our results, as LLMs are designed to handle text and not numerical values. However, for simplicity and to ease the replication or pursue of our work, we preferred to add the categorical values and the associated mapping (e.g., 1 $\rightarrow$ ”Male”, 0 $\rightarrow$ ”Female”) in the data specifications of the prompt rather than in the preprocessing step.
III-C Text-to-tabular synthetic data generation
III-C1 Approach
@@
III-C2 The text - Prompt design
@@
  • Prior knowledge: The prompt included prior knowledge of a medical database and a disease, that is, a comprehensive description of the medical database, detailing its name and nature, as well as the inclusion/exclusion criteria relevant to the population (e.g., disease, demographic criteria).
  • Instructions: We configured the LLM to minimize verbosity by instructing it not to comment on, repeat, or answer the question. Given data specifications included detailed column names, medical definitions for each variable, the type of each variable (e.g., float, integer), and the encoding of categorical variables. The output format was specified to be in JSON, with a predetermined number of rows $n$, and no missing values.
  • Context: To improve the realism of the synthetic data generated by the LLM, we leveraged its few-shot learning capabilities by including a limited number of fictitious patient examples. These examples can be derived from aggregated data found in the literature or by consulting medical experts. When using real patient records for generating examples, patient privacy must be considered. In our case, we added a single example created using the average values of the selected variables extracted from publications on #PPMI2020 [35] and on the ADNI dataset [36, 37].
The prompts designed to generate our synthetic PPMI and ADNI datasets are in the Appendix B (see Figure 3 for PPMI and Figure 4 in Supplementary Materials for ADNI).
III-C3 The tabular SDG process
@@
Ordering of columns being irrelevant in tabular datasets, we randomly permuted the columns at each generation, as proposed by [11].
III-D Evaluation framework of synthetic patient data generation
To obtain a comprehensive assessment, we evaluated both fidelity, i.e. statistical similarity, privacy preservation and utility, i.e. added value of the synthetic data to a specific ML task or healthcare use case, using several widely used metrics in the SDG field or proposed by the Synthetic Data Vault in the Python library SDMetrics [38].
III-D1 Evaluation framework
We compared our approach with several baseline tabular-to-tabular SDG Models that require access to the original data during training: CTGAN [18], TVAE [18], and Gaussain copula (GC)[16]. CTGAN and TVAE were both trained for 300 epochs and using a batch size of 50. We chose a smaller batch size than the one proposed by [18] because of the smaller training database.